OpenStudy (anonymous):

The area of a rectangle is given by the expression x2 + 5x + 4. If the length of one side is given by x + 2, what is the length of the other side?

OpenStudy (igreen):

Area of a rectangle: \(A = l \times w\) Plug in what we know: \(x^2 + 5x + 4 = l \times (x + 2)\) To find the width, we have to divide x + 2 to both sides. Can you divide this? @shelbytipple \(\dfrac{x^2+5x+4}{x+2}\)